express 0.2 M concentration in these units
1. microgram/microlitre
2. % (w/v)
3. mg/dL
Answer #1

Let us assume the molecular weight of the compound is =

M gm/mol.

Then 0.2 M solution means 1000 ml of solution contains 0.2 mole compound.

1. 0.2 mole/ litre = 0.2 * M gm / litre = 0.2 * M microgram / microlitre.

2. Weight of the compound = 0.2 * M gram

Volume = 1000 ml

So percentage of weight per volume = {(0.2*M)/1000}*100 =

0.02*M (W/V %).

3. 0.2 * M gram/ litre = 0.02* M gram/ dl = 0.02*M* 10^3 mg/dl

= 20*M mg/ dl.